OpenStudy (anonymous):

x^2+3x+8=0 find all solutions and write in a+bi form

OpenStudy (anonymous):

quadratic formula:\[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

\[\frac{-3 \pm i \sqrt{23}}{2}\] so: \[(x+\frac{3+ i \sqrt{23}}{2})(x+\frac{3-i \sqrt{23}}{2})\]